# Equilibrium in Style: A Modeling Framework on the Cash Flow and the Life Cycle of a Consumer Store

### v1

@@ 5 Part III: Potential customers
In real life, often times, we are unwilling to travel long distances to purchase from a store. This naturally suggests that willingness to pruchase is negatively correlated with distance. We set the proportion of people willing to enter the store to buy to decrease exponentially with distance. The store attracts people in the coordinate $\mathbf{x}$ at most in proportion to $e^{-\delta\|\mathbf{x}\|},$ where $\delta$ is the distance attenuation coefficient.
Moreover, at the beginning period of the store opening, there are few people that know of its existence, so it takes time to broadening its visibility. From the opening date, the store’s visibility broadens around at a speed of $k.$ Therefore, we indicate that the store will interact with a circular range less than $kt$ from the store at time $t,$ which is naturally expanding due to the broadening of the store’s visibility. Define the proportion of uncompetitive attraction as
@@
It can be seen from the proof that the continuity demonstrated by proposition 5.2 only exists in area where all stores are visible. That is to say, newly opened stores will cause an unexpected decrease in the attraction proportion of other stores during the process of broadeninging their visibility.
@@
As an explicit expression for $N_{t}$ is unavailable, Figure 8 is generated through Monte Carlo simulation (see Metropolis et al. (1953)). This limitation leads to the complexity of calibrating parameters in empirical studies, which gives motivation for us to introduce Theorem 5.1 and the competing equivalent traffic foot density.
@@
There exists a competing equivalent traffic foot density $u^{\prime}$ for a given store (say 00th store) such that
@@ 6 Part IV: Full model of cash flow
For the case where the spatial competition exists, replace $u$ with competing equivalent traffic foot density $u^{\prime}$ to get an approximation as
@@
Figure 9 shows a typical cash flow curve in the case of 10 consumer types. Here, $t=1$ is set as one day. The monetary unit of the cash flow $\text{CF}_{t}$ is Chinese yuan/ day. In Figure 9, we can see the cash flow has a rainbow-shaped curve. In the early stage, the cash flow of the store gradually increases, mainly due to the increase in the flow of potential customers, which is essentially caused by the visibility broadening. The decrease in cash flow in the later period is due to a decrease in the conversion rate caused by the style-shifting.
@@
Subsequently, we set $m=1$ and assess curve performance under varying settings: (1) population flow density $u$, (2) preference shifting parameter $\nu$, and (3) visibility broadening speed $k$. Our examination centers on two key aspects: the peak of the cash flow’s rainbow-shaped curve and the point at which it decreases by 95% from the peak (referred to as the closing point). The former represents the store’s maximum theoretical cash flow, while the latter signifies the duration of the store’s lifecycle. Figure 10 illustrates cash flow trends over time for a representative store utilizing our comprehensive model.
Above all, Figure 10 signifies the inherent lifecycle of a store has a rainbow-shaped curve. Second, a higher population flow density $u$ results in a larger peak cash flow, yet has no bearing on the store’s lifecycle. This outcome challenges conventional beliefs that a store situated in an area with low foot traffic experiences a shorter lifecycle. But our results find that each type of stores has its own viability with a response to different environments.
Third, the preference shifting parameter $\nu$ impacts both the lifecycle and the peak of the cash flow. A faster preference shift results in a shorter lifecycle and a lower peak of the cash flow. It is explained that for a store located in a shopping mall, the operational risk is larger and the ACP is also higher; Otherwise, the same investments cannot yield a similar shape.
@@ 7.1 Cash flow data and basic information of stores
This table reports the summary statistics for daily cash flow of three representative stores from three industries in China’s real economics: the retail industry (store A), the restaurant industry (store B) and the service industry (store C). Panel A displays the descriptive statistics of cash flow, while Panel B shows basic information. In particular, the ramp up period comprises of the dates that cash flow increase from zero to optimal cash flow.
@@
First of all, we collect the daily cash flow data of three stores from three representative industries of China’s real economics: the retail industry (store A), the restaurant industry (store B) and the service industry (store C)  [^2] In fact, we collects the order-level data and sum them into daily cash flow.. The data is authorized for use by the store owners. Note that the stores are usually closed on Chinese holidays (such as Lunar New Year). Therefore, we interpolate the missing value with the mean of two cash flow values in the same week day of the preceding and subsequent weeks. Table 1 displays the summary statistics for these three stores. The average cash flow of store A is 4236.45 Chinese yuan with a standard deviation of 1648.13 Chinese yuan, while that of store C is 767.75 Chinese yuan with a standard deviation of 757.89 Chinese yuan. It shows that cash flow from the service industry is more volatile than that of the retail industry, which is expected because the commodities of the retail industry is generally the same throughout and the demands of consumer is stable. However, the ACP of store A (19.24 Chinese yuan) is lower than that of store C (81.16 Chinese yuan). This result suggests that the commodity of service industry is lower frequency but higher gross profit rate.
@@
Furthermore, through real life experiences, we calibrate an important parameter: the distance attenuation coefficient $\delta$. Not many people are willing to travel more than three kilometers (km) away to make a purchase [^3] For example, see https://www.canyin88.com/zixun/2019/07/05/73882.html for more details.. Therefore, the distance attenuation coefficient $\delta$ is set as $\frac{1}{3}\ln\frac{1}{0.01}\approx 1.535$ i.e. $\hat{\delta}=1.535$, which means the attraction of the store to the foot traffic decreases to less than $1\%$ at a distance of $3$km.
@@
Estimation of traffic density by using Theorem 5.1 We collect traffic heat-maps from Baidu map, including the absolute values of the population per 100 $m^{2}$ in Yuxi City, Yunnan Province for two years, 2022 and 2023. By weighted averaging, we obtain the absolute density of the population within a three-kilometer radius of each store. In order to get a more accurate idea of the consumer density near each store, we also collect the number of comparable stores that belongs to the same category (or sub-industry) with our sample stores from Dianping, a China’s online platform where the public writes reviews about different businesses. However, the “similar style” means that the distance on the style vector of stores is sufficiently small, which does not equate the “same category”. To avoid data manipulation, we simply use the “same category” to substitute “similar styles”, a concept that defined in our paper, resulting in some problems. We will provide further discussion in 8.
@@
Therefore, we use the formula 5.1 to estimate the competitive equivalent $\hat{u}^{\prime}$. Table 2 reports the specific county, category, number of competitors, and real foot traffic density, as well as the competitive equivalent foot traffic density estimated by formula 5.1. First, store Ais in Xinping county, while store B and C are in Hongta District, showing a reasonable results that the population density (above 50000) in the district is larger than that (6246.85) in the county. Second, only two stores exist in the same category as store C (from the service industry), while 454 stores exist that are similar to store B (from the restaurant industry). The results suggest that the demand of the service industry is relatively lower than that of the rigid demand of the restaurant industry. Finally, taking the competition into account, the competitive equivalent foot traffic density of store A in the county is surprisingly highest (3470.17). This means that not only the absolute populations but also the intensity of competition matters. Overall, our paper provides a strong tool to consider both the distribution of populations and the intensity of competition.
@@
$u^{\prime}=u\frac{\delta^{2}}{2\pi}\iint_{\mathbb{R}^{2}}{e^{-\delta\|\mathbf{x}-\mathbf{x}_{0}\|-\delta d(\mathbf{x})}}\left({\sum_{j=0}^{m}e^{-\delta\|\mathbf{x}-\mathbf{x}_{j}\|}}\right)^{-1}\mathrm{d}\mathbf{x},$ where $d(\mathbf{x})=\inf_{1\leqslant j\leqslant m}\|\mathbf{x}-\mathbf{x}_{j}\|.$ In particular, we present the specific county of store A, store B, and store C as well. Furthermore, this table shows the number of competitors in the same category (or sub-industry) of the same county, where the categories is defined as the same as that in the Dianping website, the well-known China’s online platform where the public reviews stores and businesses. Furthermore, this table shows the average real foot traffic density from both 2022 and 2023 combine. The population density data stems from the Baidu map, the famous online map provider in China. Our sample period spans from January 2022 to December 2023.
@@ 7.3 Empirical results
$\text{CF}_{t}=2\pi\frac{u^{\prime}}{\delta^{2}}(1-(\delta kt+1)e^{-\delta kt})\cdot\theta(I)\tilde{\beta}_{0}(I)\exp\left(-\nu t^{2}\right)+\varepsilon_{t},$ where $\varepsilon_{t}$ is the random residual term. We set $t=1$ as one day, and the monetary unit of $\text{CF}_{t}$ is Chinese yuan/day. Furthermore, we present three representative stores distributed in three industries of China’s real economics: retail industry (model 1), restaurant industry (model 2) and service industry (model 3). Moreover, we present the estimates of $k$, $\nu$ and $\beta_{0}$ by times $10^{6}$, $10^{6}$, $10^{2}$,respectively, because the scale is too small in daily frequency. The standard error adjusted by Newey and West (1986) is enclosed in parentheses. The sample period spans from January 2022 to September 2023, during which China’s digital business rapidly grows.
@@
Panel A of Table 3 shows the NLS results of (121). First, the the decrease coefficients of store A, store B, and store C are $2.88\times 10^{-6}$ (with standard error = $0.15\times 10^{-6}$ ), $0.83\times 10^{6}$ (with standard error = $0.22\times 10^{-6}$ ) and $57.11\times 10^{-6}$ (with standard error = $12.40\times 10^{-6}$ ), respectively. It means that the decrease coefficients are all statistically significant at a 1% confidence level. The decrease speed of cash flow is shown as a exponential form, meaning that the decrease speed will increase as time goes by. For instance, it means that the cash flow of store C decreases above 1% after 68 days ( $exp(-57.11\times 10^{-6}\times 68^{2})-exp(-57.11\times 10^{-6}\times 67^{2})$). This parameter is the key factor that impacts the life cycle of stores. More importantly, the results in Panel B of Table 3 means that store C from the service industry has the longer theoretical life span (233 days), while store B from the restaurant industry has the shortest life span (1896 days). Second, store A from the retail industry has the smallest visibility broadening speed ($2.50\times 10^{-2}$ with a standard error =$0.26\times 10^{-2}$), while store B from the restaurant industry has the largest ($24.65\times 10^{-2}$ with a standard error =$15.64\times 10^{-2}$), indicating that our model performs a significant estimates at a 1% confidence level. It suggests that the people around the store knows quickly the restaurant store, possibly because of its larger monetary investment to its opening activities. Finally, the initial conversion rate of store A from retail industry is the largest (3.59% with a standard error = 0.14%), that of store C from service industry is smallest (0.17% with standard error = 0.02%), and the that of store B from restaurant industry is the middle one (3.46% with a standard error =0.01%). It shows that the parameters are all significant at 1% confidence level, corresponding with the real life experience. In our daily life, product difference among the retail store is ignored, while the service store highly relies on the service and technique of the owners. It leads to the reasonable conclusions that store Afrom retail owns the highest initial conversion rate, while store Cowns the lowest initial conversion rate.
Overall, we provide a strong empirical tools to evaluate the key parameters of the stores,including the competition-adjusted potential consumers $\hat{u}^{\prime}$, the visibility broadening speed $k$, the initial conversion rate $\beta_{0}$ and the decrease coefficient $\nu$. These parameters capture the life cycle of the stores.
@@ 7.4 Visualizations of cash flow
Figure 11 shows the raw cash flow and the fitted values of our model. Intuitively, our model perform well in capturing the trend of increase and the decrease of cash flow. More precisely, different industries have different rules in terms of frequency. First, the cash flow of the retail store C fits well in all three frequencies, suggesting that the consumer may not have special preference to one specific store, meaning that the cash flow of the retail store is nearly independently and identically distributed. However, the restaurant store B performs well in weekly frequencies, while the service store C fits well in monthly frequency. The results may indicate that these two industries have the strong seasonality. For example, we may change our taste every week, while we cut our hair monthly.
@@
Figure 11: The figure illustrate the raw cash flow and fitted values of our model on three kinds of frequency: daily, weekly and monthly. The dashed line plots the raw cash flow, while the solid one presents the fitted values of our model. The sample period spans from January 2022 to September 2023.
Overall, our model well captures the trend of cash flow of a China’s store over time by both qualitative and quantitative ways.
@@ 8.1 Some interesting implications
Our structural model provides micro-foundations for the data generation process of a store’s cash flow, from which we derive some interesting implications. First, in the real economy, though most of the stores show a rainbow-shaped cash flow, some may present as an “M” shape. Our model implies that this is due to the diversity of consumer preferences. In the beginning, the style of the store may be outdated for some consumers, but may be ahead of time for others, which means cash flow may experience increases and decreases. Second, Figure 10 shows that faster broadening of visibility not only shortens the ramp-up period but also increases the height of peak cash flow. This means that we should encourage the owner to increase monetary investment into opening promotion activities. Third, the decrease coefficient $\nu$ is the key parameter to determine the life cycle of a store. To avoid the shifting of consumer preference, the stores should frequently update their products to keep up with style. Stores that have higher budgets for product innovation will have an advantage in style updating and thus their life cycle will be longer. Similarly, policymakers should encourage the development of new product research departments. For example, a Chinese milk tea store, NAIXUE, listed in the Hong Kong exchange market update their products per month.

### v2

@@ 5 Part III: Potential customers
In real life, oftentimes, we are unwilling to travel long distances to purchase from a store. This naturally suggests that willingness to purchase is negatively correlated with distance. We set the proportion of people willing to enter the store to buy to decrease exponentially with distance. The store attracts people in the coordinate $\mathbf{x}$ at most in proportion to $e^{-\delta\|\mathbf{x}\|},$ where $\delta$ is the distance attenuation coefficient.
Moreover, at the beginning period of the store opening, there are few people who know of its existence, so it takes time to broaden its visibility. From the opening date, the store’s visibility broadens around at a speed of $k.$ Therefore, we indicate that the store will interact with a circular range less than $kt$ from the store at time $t,$ which is naturally expanding due to the broadening of the store’s visibility. Define the proportion of uncompetitive attraction as
@@
It can be seen from the proof that the continuity demonstrated by proposition 5.2 only exists in area where all stores are visible. That is to say, newly opened stores will cause an unexpected decrease in the attraction proportion of other stores during the process of broadening their visibility.
@@
As an explicit expression for $N_{t}$ is unavailable, Figure 8 is generated through Monte Carlo simulation (see Metropolis et al. (1953)). This limitation leads to the complexity of calibrating parameters in empirical studies, which gives motivation for us to introduce Theorem 5.1 and the competing equivalent foot traffic density.
@@
There exists a competing equivalent foot traffic density $u^{\prime}$ for a given store (say 00th store) such that
@@ 6 Part IV: Full model of cash flow
For the case where the spatial competition exists, replace $u$ with competing equivalent foot traffic density $u^{\prime}$ to get an approximation as
@@
Figure 9 shows a typical cash flow curve in the case of 10 consumer types. Here, $t=1$ is set as one day. The monetary unit of the cash flow $\text{CF}_{t}$ is Chinese yuan/ day. In Figure 9, we can see the cash flow has a rainbow-shaped curve. In the early stage, the cash flow of the store gradually increases, mainly due to the increase in the flow of potential customers, which is essentially caused by the visibility broadening. The decrease in cash flow in the later period is due to a decrease in the conversion rate caused by style-shifting.
@@
Subsequently, we set $m=1$ and assess curve performance under varying settings: (1) population flow density $u$, (2) decrease coefficient $\nu$, and (3) visibility broadening speed $k$. Our examination centers on two key aspects: the peak of the cash flow’s rainbow-shaped curve and the point at which it decreases by 95% from the peak (referred to as the closing point). The former represents the store’s maximum theoretical cash flow, while the latter signifies the duration of the store’s lifecycle. Figure 10 illustrates cash flow trends over time for a representative store utilizing our comprehensive model.
Above all, Figure 10 signifies the inherent lifecycle of a store has a rainbow-shaped curve. Second, a higher foot traffic density $u$ results in a larger peak cash flow yet has no bearing on the store’s lifecycle. This outcome challenges conventional beliefs that a store situated in an area with low foot traffic experiences a shorter lifecycle. But our results find that each type of stores has its own viability with a response to different environments.
Third, the decrease coefficient $\nu$ impacts both the lifecycle and the peak of the cash flow. A faster preference shift results in a shorter lifecycle and a lower peak of the cash flow. It is explained that for a store located in a shopping mall, the operational risk is larger and the ACP is also higher; Otherwise, the same investments cannot yield a similar shape.
@@ 7.1 Cash flow data and basic information of stores
This table reports the summary statistics for the daily cash flow of three representative stores from three industries in China’s real economics: the retail industry (store A), the restaurant industry (store B), and the service industry (store C). Panel A displays the descriptive statistics of cash flow, while Panel B shows basic information. In particular, the ramp-up period comprises the dates that cash flow increases from zero to the optimal cash flow.
@@
First of all, we collect the daily cash flow data of three stores from three representative industries of China’s real economics: the retail industry (store A), the restaurant industry (store B) and the service industry (store C)  [^2] In fact, we collects the order-level data and sum them into daily cash flow.. The data is authorized for use by the store owners. Note that the stores are usually closed on Chinese holidays (such as Lunar New Year). Therefore, we interpolate the missing value with the mean of two cash flow values on the same weekday of the preceding and subsequent weeks. Table 1 displays the summary statistics for these three stores. The average cash flow of store A is 4236.45 Chinese yuan with a standard deviation of 1648.13 Chinese yuan, while that of store C is 767.75 Chinese yuan with a standard deviation of 757.89 Chinese yuan. It shows that cash flow from the service industry is more volatile than that of the retail industry, which is expected because the commodities of the retail industry are generally the same throughout and the demands of consumers are stable. However, the ACP of store A (19.24 Chinese yuan) is lower than that of store C (81.16 Chinese yuan). This result suggests that the commodity of the service industry is lower frequency but higher gross profit rate.
@@
Furthermore, through real-life experiences, we calibrate an important parameter: the distance attenuation coefficient $\delta$. Not many people are willing to travel more than three kilometers (km) away to make a purchase [^3] For example, see https://www.canyin88.com/zixun/2019/07/05/73882.html for more details.. Therefore, the distance attenuation coefficient $\delta$ is set as $\frac{1}{3}\ln\frac{1}{0.01}\approx 1.535$ i.e. $\hat{\delta}=1.535$, which means the attraction of the store to the foot traffic decreases to less than $1\%$ at a distance of $3$km.
@@
Estimation of traffic density by using Theorem 5.1 We collect traffic heat-maps from Baidu map, including the absolute values of the population per 100 $m^{2}$ in Yuxi City, Yunnan Province for two years, 2022 and 2023. By weighted averaging, we obtain the absolute density of the population within a three-kilometer radius of each store. In order to get a more accurate idea of the consumer density near each store, we also collect the number of comparable stores that belong to the same category (or sub-industry) with our sample stores from Dianping, a China’s online platform where the public writes reviews about different businesses. However, the “similar style” means that the distance on the style vector of stores is sufficiently small, which does not equate to the “same category”. To avoid data manipulation, we simply use the “same category” to substitute “similar styles”, a concept that is defined in our paper, resulting in some problems. We will provide further discussion in 8.
@@
Therefore, we use the formula 5.1 to estimate the competitive equivalent $\hat{u}^{\prime}$. Table 2 reports the specific county, category, number of competitors, and real foot traffic density, as well as the competitive equivalent foot traffic density estimated by the formula (5.1). First, store A is in Xinping County, while stores B and C are in Hongta District, showing reasonable results that the population density (above 50000) in the district is larger than that (6246.85) in the county. Second, only two stores exist in the same category as store C (from the service industry), while 454 stores exist that are similar to store B (from the restaurant industry). The results suggest that the demand of the service industry is relatively lower than that of the rigid demand of the restaurant industry. Finally, taking the competition into account, the competitive equivalent foot traffic density of store A in the county is surprisingly highest (3470.17). This means that not only the absolute population but also the intensity of competition matters. Overall, our paper provides a strong tool to consider both the distribution of populations and the intensity of competition.
@@
$u^{\prime}=u\frac{\delta^{2}}{2\pi}\iint_{\mathbb{R}^{2}}{e^{-\delta\|\mathbf{x}-\mathbf{x}_{0}\|-\delta d(\mathbf{x})}}\left({\sum_{j=0}^{m}e^{-\delta\|\mathbf{x}-\mathbf{x}_{j}\|}}\right)^{-1}\mathrm{d}\mathbf{x},$ where $d(\mathbf{x})=\inf_{1\leqslant j\leqslant m}\|\mathbf{x}-\mathbf{x}_{j}\|.$ In particular, we present the specific county of store A, store B, and store C as well. Furthermore, this table shows the number of competitors in the same category (or sub-industry) of the same county, where the categories are defined as the same as that in the Dianping website, the well-known China’s online platform where the public reviews stores and businesses. Furthermore, this table shows the average real foot traffic density from both 2022 and 2023 combined. The population density data stems from the Baidu map, the famous online map provider in China. Our sample period spans from January 2022 to December 2023.
@@ 7.3 Empirical results
$\text{CF}_{t}=2\pi\frac{u^{\prime}}{\delta^{2}}(1-(\delta kt+1)e^{-\delta kt})\cdot\theta(I)\tilde{\beta}_{0}(I)\exp\left(-\nu t^{2}\right)+\varepsilon_{t},$ where $\varepsilon_{t}$ is the random residual term. We set $t=1$ as one day, and the monetary unit of $\text{CF}_{t}$ is Chinese yuan/day. Furthermore, we present three representative stores distributed in three industries of China’s real economics: the retail industry (model 1), the restaurant industry (model 2), and the service industry (model 3). Moreover, we present the estimates of $k$, $\nu$, and $\beta_{0}$ by times $10^{6}$, $10^{6}$, $10^{2}$, respectively, because the scale is too small in daily frequency. The standard error adjusted by Newey and West (1986) is enclosed in parentheses. The sample period spans from January 2022 to September 2023, during which China’s digital business rapidly grows.
@@
Panel A of Table 3 shows the NLS results of (121). First, the decrease coefficients of store A, store B, and store C are $2.88\times 10^{-6}$ (with standard error = $0.15\times 10^{-6}$ ), $0.83\times 10^{6}$ (with standard error = $0.22\times 10^{-6}$ ) and $57.11\times 10^{-6}$ (with standard error = $12.40\times 10^{-6}$ ), respectively. It means that the decrease coefficients are all statistically significant at a 1% confidence level. The decreased speed of cash flow is shown as an exponential form, meaning that the decreased speed will increase as time goes by. For instance, it means that the cash flow of store C decreases above 1% after 68 days ( $exp(-57.11\times 10^{-6}\times 68^{2})-exp(-57.11\times 10^{-6}\times 67^{2})$). This parameter is the key factor that impacts the life cycle of stores. More importantly, the results in Panel B of Table 3 means that store C from the service industry has the longer theoretical life span (233 days), while store B from the restaurant industry has the shortest life span (1896 days). Second, store A from the retail industry has the smallest visibility broadening speed ($2.50\times 10^{-2}$ with a standard error =$0.26\times 10^{-2}$), while store B from the restaurant industry has the largest ($24.65\times 10^{-2}$ with a standard error =$15.64\times 10^{-2}$), indicating that our model performs a significant estimate at a 1% confidence level. It suggests that the people around the store know quickly the restaurant store, possibly because of its larger monetary investment in its opening activities. Finally, the initial conversion rate of store A from the retail industry is the largest (3.59% with a standard error = 0.14%), that of store C from the service industry is the smallest (0.17% with standard error = 0.02%), and that of store B from restaurant industry is the middle one (3.46% with a standard error =0.01%). It shows that the parameters are all significant at 1% confidence level, corresponding with the real-life experience. In our daily lives, product difference among retail stores is ignored, while the service store highly relies on the service and technique of the owners. It leads to the reasonable conclusion that store A from retail owns the highest initial conversion rate, while store C owns the lowest initial conversion rate.
Overall, we provide a strong empirical tool to evaluate the key parameters of the stores, including the competition-adjusted potential consumers $\hat{u}^{\prime}$, the visibility broadening speed $k$, the initial conversion rate $\beta_{0}$ and the decrease coefficient $\nu$. These parameters capture the life cycle of the stores.
@@ 7.4 Visualizations of cash flow
Figure 11 shows the raw cash flow and the fitted values of our model. Intuitively, our model performs well in capturing the trend of increase and decrease in cash flow. More precisely, different industries have different rules in terms of frequency. First, the cash flow of the retail store C fits well in all three frequencies, suggesting that the consumer may not have a special preference for one specific store, meaning that the cash flow of the retail store is nearly independently and identically distributed. However, the restaurant store B performs well in weekly frequencies, while the service store C fits well in monthly frequencies. The results may indicate that these two industries have a strong seasonality. For example, we may change our taste every week, while we cut our hair monthly.
@@
Figure 11: The figure illustrates the raw cash flow and fitted values of our model on three kinds of frequency: daily, weekly and monthly. The dashed line plots the raw cash flow, while the solid one presents the fitted values of our model. The sample period spans from January 2022 to September 2023.
Overall, our model well captures the trend of cash flow of China’s stores over time in both qualitative and quantitative ways.
@@ 8.1 Some interesting implications
Our structural model provides micro-foundations for the data generation process of a store’s cash flow, from which we derive some interesting implications. First, in the real economy, though most of the stores show a rainbow-shaped cash flow, some may present as an “M” shape. Our model implies that this is due to the diversity of consumer preferences. In the beginning, the style of the store may be outdated for some consumers but may be ahead of time for others, which means cash flow may experience increases and decreases. Second, Figure 10 shows that faster broadening of visibility not only shortens the ramp-up period but also increases the height of peak cash flow. This means that we should encourage the owner to increase monetary investment into opening promotion activities. Third, the decrease coefficient $\nu$ is the key parameter to determine the life cycle of a store. To avoid the shifting of consumer preference, the stores should frequently update their products to keep up with style. Stores that have higher budgets for product innovation will have an advantage in style updating and thus their life cycle will be longer. Similarly, policymakers should encourage the development of new product research departments. For example, a Chinese milk tea store, NAIXUE, listed in the Hong Kong exchange market updates its products per month.